(15 points) Triangular Arbitrage w/ bid-ask spreads Continue to assume that you are a large money...

  1. (15 points) Triangular Arbitrage w/ bid-ask spreads Continue to assume that you are a large money center bank and you have 3,000,000 South Korean Won. The following banks make the following quotes:

Bank

Location of Bank

Exchange rate quotes

Bank BNP Paribas

Paris, France

137.1472 yen/euro
137.1463 yen/euro

The Bank of Tokyo - Mitsubishi UFJ

Tokyo, Japan

10.1251 won/yen
10.1247 won/yen

Korea Exchange Bank

Seoul, Korea

1386.57 won/euro
1386.42 won/euro

Determine a way to make a profit. Using complete sentences, describe the actions you would take to make this profit, stating the amounts of currency you would exchange in each step and through what institution you make that exchange. Also, clearly state the profit you make and be clear what currency that profit is in.
Answer #1

We have 3,000,000 won at the beginning

Now, we will visit Korea Exchange Bank and exchange 3,000,000 won for euro

The exchange rate used will be the bid rate i.e. 1386.57 won/euro as we are buying the base currency

Thus, we will receive (3,000,000/1386.57) = 2163.6124 euros

Now, we will visit Bank BNP Paribas and convert the euro into yen

We will use ask rate i.e. 137.1463 yen/euro for this transaction as we are selling the base currency

Thus, we will receive (2163.6124 * 137.1463) = 296,731.4308 yens

Now, we will visit The Bank of Tokyo and exchange the yens for won at the ask rate i.e 10.1247 won/yen as we are selling the base currency

So we will receive, (296,731.4308 * 10.1247) = 3,004,316.7174 won

Thus, we ended up with 3,004,316.7174 won which means we made a profit of 3,004,316.7174 - 3,000,000 won = 4,316.7174 wons